# Deep Learning Based Dynamics Identification and Linearization of Orbital Problems using Koopman Theory

### v2

@@ 3.1 Deep Neural Network Structure and Loss Function Definition
In order to further improve the accuracy of the model learned, training techniques such as $L_{1}$ and $L_{2}$ regularization are implemented into the loss function as a means to penalize unwanted behaviours by the DNN [59]. For instance, $L_{1}$ regularization aims to penalize the model for having large weights, achieving this objective by including the sum of the absolute value of the weights in the loss function, thus calculated as:
@@
$L_{2}$ regularization on the other hand, encourages the sum of the squares of the weights to be small, thus reducing the chance of overfitting, whilst improving the networks ability to learn complex features. This is particularly useful when handling nonlinear dynamics of increased dimension and increased terms as in the CR3BP. The $L_{2}$ regularisation loss function can be calculated as:
@@
$\mathcal{L}_{\text{total}}=\gamma\mathcal{L}_{\text{Pred}}+\beta\mathcal{L}_{\text{Recon}}+\lambda_{1}\mathcal{L}_{L_{1}}+\lambda_{2}\mathcal{L}_{L_{2}}$ (32)
The total loss function value, $\mathcal{L}_{\text{total}}$ is then used in the Adam optimizer, to recursively improve the weights and biases of the DNN, so that the observable functions can be appropriately learned. With improved observable functions, the approximation to the Koopman operator becomes more accurate, hence the performance of the linear system becomes more true. The weights $\gamma,\beta,\lambda_{1},\lambda_{2}$ corresponding to each loss function, can be adjusted according to the performance of the network and knowledge of the dynamics of the system.
@@ 4.1 Two-Body Circular Metrics
    $r\equiv|{\bf r}|=\sqrt{x^{2}+y^{2}}$
@@
    $v\equiv|{\bf v}|=\sqrt{\dot{x}^{2}+\dot{y}^{2}}$
@@
    ${\bf r}\cdot{\bf v}=x\dot{x}+y\dot{y}$
@@
    $L_{z}\equiv\left({\bf r}\times{\bf v}\right)_{z}=x\dot{y}-y\dot{x}$
@@
From the determination of the third metric (namely #3), which is ${\bf r}\cdot{\bf v}$, in the circular 2BP, it is found that the numerical value of this metric follows a sinusoidal shape for one orbital period of calculation. Hence, when the average of the entire dataset is taken, this value tends toward zero, consequently – when applied to the relative error formula given by (36) – yielding a substantially greater magnitude of error than what is actually observed in the simulation. Therefore, in the discussion of the ${\bf r}\cdot{\bf v}$ metric, it should be noted that the metric is not a relative error but a plot of the actual value. We can then determine the efficacy of this metric by determining how far the value is from zero at every calculated point.
@@ 4.2 CR3BP Jacobi Constant
In order to determine the accuracy of the algorithm on predicting a true representation of the CR3BP, the aforementioned metrics would not suffice, because they are exclusively applicable to the circular 2BP. We can however, utilise the conserved quantity known as the Jacobi Constant, which is a relative negative measure of energy [9, 61] to ensure whether or not our method is satisfactory. Given that the Jacobi constant is conserved, and it is the only constant of motion in the CR3BP (since it has no ignorable coordinates [62]), hence it is a metric that we can employ on our predicted Koopman dynamics. The rigorous proof of the Jacobi constant can be derived in numerous ways [9, 63, 57], however the derivation of the equation used in this work mirrors [64], with the Jacobi constant defined as:
@@ 5 SIMULATION AND RESULTS
In this section, we present the simulation, results and discussion of the implementation of our proposed RLDK method. A comparison against the nonlinear dynamics, and an investigation of the aforementioned metrics for both the 2BP and the CR3BP are carried out. The full PyTorch code for data generation, training and the examples provided can be found on GitHub [^1] https://github.com/tiwari-research-group/Koopman-Orbital-Dynamics.
@@ 5.1 Data Generation
In order to adequately train the DNN, a large dataset of varying initial conditions (IC’s) associated with orbital propagation is collected prior. The method in which we derive a range of different IC’s for the training data is slightly different for both the 2BP and CR3BP, however the data generation method after the IC has been selected is the same for both dynamics – refer to Algorithm 2. In the training of the DNN, a training-validation-test split of $80\%-10\%-10\%$ was utilized.
@@
5:     Generate random I.C based on 23 or 28
@@
5.1.1 Circular Two-Body Data Generation
In order to create a 2BP circular orbit, the correct IC’s must be selected. The semi-major axis $a$ and eccentricity $e$ or an orbit are the only orbital elements needed to parameterize the shape of the orbit. Hence, only these will be calculated and used to determine the IC for the particular orbit. In particular the orbital element which will be randomly varied in the generation of the data is the semi-major axis $a$. This is done implicitly by randomly choosing an altitude between 200km and 5000km and using the following relation:
@@ 5.1.1 Circular Two-Body Data Generation
where $R_{E}=6378.14km$ is the radius of the Earth (assumed to be spherical) and $h$ is the randomly chosen altitude. Naturally, because the orbits we wish to generate are circular, we set the eccentricity to zero, $e=0$. From this information, we can use equations (22) and (23) to determine the initial positions and velocities for the current orbit. Each orbit is propagated for the length of one period ($T=2\pi\sqrt{\frac{a^{3}}{\mu}}$) before being truncated by $\alpha$ data points to accommodate the prediction loss function outlined in (29). For the 2BP, $\alpha=25$. The number of IC’s in each training set was 100 for the 2BP.
@@ 5.2.1 Two-Body Problem Linearization
The hyperparameters of the DNN used for all experiments with the 2BP are outlined in Table 1. Total training time for the DNN used in the 2BP was 47 minutes on an NVIDIA GeForce RTX 3090 GPU.
@@
Figure 2 shows the prediction by the approximated Koopman operator on a range of orbits of varying semi-major axes. The green shaded region represents the range of values that are used for training. It can be noted that the prediction of the orbits are very accurate for regions both inside and outside the training region, aligning with the true dynamics very closely. The Koopman operator is able to globally linearize the trajectory for any altitude that a satellite in orbit around the Earth would be placed into. Figure 3 displays the four states of the satellite in each varying altitude trajectory and shows that not only position, but the velocity of the satellite is modelled quite well. It is clear however, that as the orbit progresses, the prediction error in the position states is somewhat periodic, but not perfectly sinusoidal as seen in Figure 4.
Figure 2: Koopman operator prediction and comparison against nonlinear dynamics of 5 orbits around Earth at various altitudes.
@@
Figure 4: Error in position states over one orbital period for Earth.
The Koopman operator, although excellent at predicting learned nonlinear dynamics, can often struggle when generalized and implemented in other systems. Here, we show that the model proposed in this work generalizes quite accurately to two other 2BP’s of varying sizes. In particular, orbital trajectories around the Moon (for a smaller central body) and Jupiter (for a larger central body) were predicted using the the same Koopman operator learned from the dataset based on Earth’s orbit. As seen in Figures 7 and 8, the error in both the Moon’s and Jupiter’s orbits are slightly higher than that of the Earth’s, but the prediction still is highly effective. Figures 5 and 6 display the prediction of orbits for the previously mentioned set of $5$ different altitudes.
Figure 5: Koopman operator prediction and comparison against nonlinear dynamics of 5 orbits around the Moon at various altitudes.
Figure 6: Koopman operator prediction and comparison against nonlinear dynamics of 5 orbits around Jupiter at various altitudes.
Plots of the error in the position state for both the Moon- and Jupiter-centered orbits allow us to analyse the effect of the size of the central body on the models ability to evolve the dynamics. A larger central body appears to have a greater periodic prediction error as compared to a central body that is smaller than that of the body the training data was derived from. As such, the possibility for the model to be used as a baseline, with additional data from the applied system added online, could enhance the performance of the model. With a short training time of around 80 min, it is not unreasonable to engage in the possibility of online data collection and learning to further enhance the accuracy of the predicted orbit.
@@
Figure 7: Error in position states over one orbital period for the Moon.
@@
Figure 8: Error in position states over one orbital period for Jupiter.
@@ 5.2.2 Accuracy Metrics
We now look at the metrics provided in Section 4.1, and how the proposed Koopman operator is able to adhere to the conservative properties aforementioned. We see in Figure 9 that the values of all four metrics for the model around Earth are substantially low and very close to zero, hence illustrating the absolute error in each metric along the orbit is near zero. Furthermore, when these metrics are checked on orbits around both the Moon and Jupiter, Figures 10 and 11, the conservation of the invariant properties is also preserved, as the values are near zero. These metrics prove the accuracy and reliability of our model in representing the 2BP not only insofar as its specific trained dataset is concerned, but also in its generalized applications to other systems.
@@ 5.2.3 Circular Restricted Three-Body Problem Linearization
Figure 13 shows the linearized evolution of the dynamics that is generated by our Koopman operator. It shows that the DNN is able to learn appropriate observables and lift the dimension of the states such that the K matrix was able to accurately predict the orbit of the CR3BP by rendering it as a linear system. Because the orbit, like the 2BP, is in-plane, the $z$-axis position and velocity are zero, therefore we are able to plot the propagation of the dynamics on a 2D plane to further inspect the accuracy of the propagation, as done in Figure 14. Note that in both Figures 13 and 14 the scale of the Earth and Moon representations are chosen for ease of viewing, and are not to scale with the propagated orbit. We also show the evolution of each of the 4 states independently in Figure 12, comparing them to the nonlinear dynamics, to emphasise the ability to closely follow the intricate dynamics of the CR3BP. It shall be noted that although the CR3BP is propagated in the non-dimensionalized units, in these figures the dimensionalized units are restored, and are displayed in the rotating frame.

### v3

@@ 3.1 Deep Neural Network Structure and Loss Function Definition
In order to further improve the accuracy of the model learned, training techniques such as $L_{1}$ and $L_{2}$ regularization are implemented into the loss function as a means to penalize unwanted behaviors by the DNN [61]. For instance, $L_{1}$ regularization aims to penalize the model for having large weights, achieving this objective by including the sum of the absolute value of the weights in the loss function, thus calculated as:
@@
$L_{2}$ regularization on the other hand, encourages the sum of the squares of the weights to be small, thus reducing the chance of overfitting, whilst improving the networks ability to learn complex features. This is particularly useful when handling nonlinear dynamics of increased dimension and increased terms as in the CR3BP. The $L_{2}$ regularization loss function can be calculated as:
@@
Lastly, because we have knowledge of the system that we desire to represent, we can make use of some physics-informed loss functions to further help our network model the physical properties of the system as accurately as possible. Section 4.2 explains the accuracy metrics that are used to ensure physical consistency between our model and the exact nonlinear dynamics however for the 2BP model, we utilize the third metric, $r\cdot v$, as an additional loss function. We again use the MSE formula to generate the loss, and because we desire this metric to be constantly zero, we use that as our truth to remove numerical errors:
$\mathcal{L}_{r\cdot v}=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}\left\|{\hat{(r\cdot v)}-0}\right\|_{2}^{2}$ (33)
@@
$\mathcal{L}_{\text{total}}=\gamma\mathcal{L}_{\text{Pred}}+\beta\mathcal{L}_{\text{Recon}}+\lambda_{1}\mathcal{L}_{L_{1}}+\lambda_{2}\mathcal{L}_{L_{2}}+\lambda_{r\cdot v}\mathcal{L}_{r\cdot v}$ (34)
The total loss function value, $\mathcal{L}_{\text{total}}$ is then used in the Adam optimizer, to recursively improve the weights and biases of the DNN, so that the observable functions can be appropriately learned. With improved observable functions, the approximation to the Koopman operator becomes more accurate, hence the performance of the linear system becomes more true. The weights $\gamma,\beta,\lambda_{1},\lambda_{2},\lambda_{r\cdot v}$ corresponding to each loss function, can be adjusted according to the performance of the network and knowledge of the dynamics of the system.
@@ 4 MODEL ACCURACY METRICS
4.1 Global vs Local Error
In dynamic prediction there is a distinction between the global error in prediction and the local error. While the local error, $e_{n}$, only considers the error between the predicted model and the exact model over one time-step, the global error, $E_{n}$ considers the error over all time-steps and is a more accurate measure of model performance [62]. An important distinction is the fact that the global error is not equivalent to the summation of all local errors, this is shown in [62]. The definitions for both errors are given below, and we highlight our model performance by displaying both errors for all simulations in the 2BP problem.
$e_{n}\equiv\mathbf{\Phi}(\boldsymbol{x}_{n})-\boldsymbol{K}\mathbf{\Phi}(\boldsymbol{x}_{n-1})$ (35)
$E_{n}\equiv\mathbf{\Phi}(\boldsymbol{x}_{n})-\boldsymbol{K^{n}}\mathbf{\Phi}(\boldsymbol{x}_{0})\equiv\mathbf{\Phi}(\boldsymbol{x}_{n})-\boldsymbol{K}\hat{\mathbf{\Phi}}(\boldsymbol{x}_{n-1})$ (36)
where the recursive nature of the estimated state $\hat{\mathbf{\Phi}}$ is defined in Equation 16.
@@ 4.2 Two-Body Circular Metrics
    $r\equiv|{\bf r}|=\sqrt{x^{2}+y^{2}}$ The orbital radius, $r$, of a circular orbit should remain constant throughout the entire orbit. This constitutes the definition of a circular orbit.
@@
    $v\equiv|{\bf v}|=\sqrt{\dot{x}^{2}+\dot{y}^{2}}$ The magnitude of velocity, $v$ of a circular orbit should remain constant throughout the entire orbit.
@@
    ${\bf r}\cdot{\bf v}=x\dot{x}+y\dot{y}$ The velocity vector in a circular orbit should remain perpendicular to the radial direction, hence their dot product should be zero.
@@
    $L_{z}\equiv\left({\bf r}\times{\bf v}\right)_{z}=x\dot{y}-y\dot{x}$ The angular momentum, $L_{z}$, of a circular orbit, defined as the cross product of the orbital radius with the tangential velocity should be a constant value.
@@
From the determination of the third metric (namely #3), which is ${\bf r}\cdot{\bf v}$, in the circular 2BP this value should be equivalent to zero as the velocity vector should always be tangential to the positional vector, however in numerical simulation it is found that the value of this metric follows a small sinusoidal shape for one orbital period of calculation. Hence, when the average of the entire dataset is taken, this value tends toward zero, consequently – when applied to the relative error formula given by (40) – yields a substantially greater magnitude of error than what is actually observed in the simulation. Therefore, in the discussion of the ${\bf r}\cdot{\bf v}$ metric, it should be noted that the metric is not a relative error but a plot of the actual value. We can then determine the efficacy of this metric by determining how far the value is from zero at every calculated point.
@@ 4.3 CR3BP Jacobi Constant
In order to determine the accuracy of the algorithm on predicting a true representation of the CR3BP, the aforementioned metrics would not suffice, because they are exclusively applicable to the circular 2BP. We can however, utilize the conserved quantity known as the Jacobi Constant, which is a relative negative measure of energy [9, 64] to ensure whether or not our method is satisfactory. Given that the Jacobi constant is conserved, and it is the only constant of motion in the CR3BP (since it has no ignorable coordinates [65]), hence it is a metric that we can employ on our predicted Koopman dynamics. The rigorous proof of the Jacobi constant can be derived in numerous ways [9, 66, 59], however the derivation of the equation used in this work mirrors [67], with the Jacobi constant defined as:
@@ 5 SIMULATION AND RESULTS
In this section, we present the simulation, results and discussion of the implementation of our proposed method. We perform a simulations of the pure circular and elliptical nonlinear dynamics of the 2BP, the SRP and J2 perturbed circular 2BP and the CR3BP around the L1 Lagrange point. We also perform numerical analysis of the aforementioned physical accuracy metrics for both the 2BP and the CR3BP. The full PyTorch code for data generation, training and the examples provided can be found on GitHub [^1] https://github.com/tiwari-research-group/Orbital-Koopman.
@@ 5.1 Data Generation
In order to adequately train the DNN, a large dataset of varying initial conditions (IC’s) associated with orbital propagation is collected prior. The method in which we derive a range of different IC’s for the training data is slightly different for both the variations of the 2BP and CR3BP, however the data generation method after the IC has been selected is the same for both dynamics – refer to Algorithm 2.
@@
5:     Generate random I.C based on 24 or 29
@@
5.1.1 Two-Body Data Generation
In order to create either the 2BP circular or elliptical orbit, the correct IC’s must be selected. The semi-major axis $a$ and eccentricity $e$ of an orbit are the only orbital elements needed to parameterize the shape of the orbit. Hence, only these will be calculated and used to determine the IC for the particular orbit. In particular the orbital element which will be randomly varied in the generation of the data is the semi-major axis $a$. This is done implicitly by randomly choosing an altitude between 200 km and 5000 km and using the following relation:
@@ 5.1.1 Two-Body Data Generation
where $R_{E}=6378.14km$ is the radius of the Earth (assumed to be spherical) and $h$ is the randomly chosen altitude. For the pure circular and perturbed circular orbits, we set $e=0$ whilst for the elliptical orbits the eccentricity is sampled from the range $[0.1-0.5]$. From this information, we can use equations (23) and (24) to determine the initial positions and velocities for the current orbit. Each orbit is propagated for the length of one period ($T=2\pi\sqrt{\frac{a^{3}}{\mu}}$) before being truncated by $\alpha$ data points to accommodate the prediction loss function outlined in (30). For the 2BP, $\alpha=25$. The number of IC’s in each training set was 200 for the 2BP.
The generation of SRP in the perturbed orbits requires a Sun position vector, from which the direction of perturbation can be calculated. As the training data contains only one period of an orbit, it is short enough to assume that the position vector of the Sun is constant. We take the position of the Sun to be in the negative X-direction (i.e. $[-1,0]$).
@@ 5.2.1 Two-Body Problem Linearization
The hyperparameters of the DNN used for all experiments with the 2BP are outlined in Table 1. Total training time for the DNN used in the 2BP was 2 hours on an NVIDIA GeForce RTX 3090 GPU. The hyperparameters of both networks are tuned emperically, whilst it is noted that large networks (both in depth and width) do not necessarily result in improved results.
@@
Figure 2 shows the prediction by the approximated Koopman operator on a range of purely circular orbits of varying semi-major axes. It can be noted that the prediction of the orbits are very accurate for regions both inside and outside the training region, aligning with the nonlinear dynamics very closely. Figures 3 and 4 display the global and local errors in prediction as outlined in Section 4.1. The largest global positional errors in the single orbit prediction are on the order of 2 km, equivalent to $0.018\%$ of the radius of the orbit, whilst the local error has an maximum positional 8 m, around $2.16e^{-6}\%$ of the orbital radius. It is evident from Figure 2 that for each varying altitude trajectory not only is position, but the velocity of the satellite is also modeled quite well.
@@
Figure 3: Global error in position states over one orbital period around Earth.
@@
Figure 4: Local error in position states over one orbital period around Earth.
It is clear however, that as the orbit progresses, the prediction error in the position states is somewhat periodic, but not perfectly sinusoidal and appears to grow as the prediction increases. As such, the following simulation demonstrates the prediction of the linear Koopman model over a length of 10 orbital periods, a length of time 10 times greater than any of the training data the model was exposed to. Figure 5 displays the four states over the entire simulation duration, where it is evident that the linear Koopman model closely follows the nonlinear dynamics. Again, the global and local errors are presented to thoroughly analyze the ability of the model to closely predict the states of system with Figures 6 and 7. We can see that although the global error slowly grows as the prediction time evolves, the local error is relatively unchanged regardless of the time length as expected. We note that the maximum global positional error is only $1.8\%$ of the orbital radius.
Figure 5: Four position and velocity states of the Earth orbits over ten orbital periods in comparison to the full nonlinear dynamics.
@@
Figure 6: Global error in position states over ten orbital periods around Earth.
Figure 7: Local error in position states over ten orbital periods around Earth.
The Koopman operator, although excellent at predicting learned nonlinear dynamics, can often struggle when generalized and implemented in other systems. Here, we show that the model proposed in this work generalizes quite accurately to two other 2BP’s of varying sizes. In particular, orbital trajectories around the Moon (for a smaller central body) and Jupiter (for a larger central body) were predicted using the same Koopman operator learned from the dataset based on Earth’s orbit. As seen in Figures [8- 11], the global and local errors in both the Moon’s and Jupiter’s orbits are slightly higher than that of the Earth’s, but the prediction still is highly effective.
Figure 8: Global error in position states over one orbital period for the Moon.
Figure 9: Local error in position states over one orbital period for the Moon.
Figure 10: Global error in position states over one orbital period for Jupiter.
Figure 11: Local error in position states over one orbital period for Jupiter.
In order to truly analyze the generality of our model to capture many of the key aspects of the real 2BP this next simulation presents the scenario where a circular orbit around Earth is perturbed due to solar radiation pressure and the J2 gravity model. Utilizing a training dataset which includes these perturbations, we are able to use our same framework with only a modification in the number of neurons per hidden layer from 25 to 35, to achieve reasonably accurate results. The predicted states of the system show close correspondence to the nonlinear system as shown in Figure 12, whilst the global and local errors in Figures 13 and 14 remain within $1.3\%$ of the orbital radius. It can be noted that further tuning and adjustments to the hyperparameters of the network would aid the ability of the model to account for these and other perturbations that are present in realistic simulations.
Figure 12: Four position and velocity states of the J2 and SRP perturbed orbits over one orbital period in comparison to the full nonlinear dynamics.
Figure 13: Global error in position over one orbital period of the J2 and SRP perturbed orbit.
Figure 14: Local error in position over one orbital period of the J2 and SRP perturbed orbit.
The next set of simulations come from modeling an eccentric 2BP where the eccentricities of the orbits are 0.1, 0.2, and 0.5. For this simulation, nonlinear orbits with varying eccentricities were included in the training dataset to enable the model to learn this particular system. The states of the various eccentric orbits are presented in Figure 15, whilst the global and local errors, as shown previously are given in Figures 16 and 17. Whilst the errors in position are relatively low and comparable to that of the purely circular or perturbed circular orbits, it is evident from Figure 15 that for orbits with increasing eccentricities, the velocity prediction begins to degrade under the current conditions.
Figure 15: States of the eccentric Earth orbits over one orbital period in comparison to the full nonlinear dynamics.
Figure 16: Global error in eccentric position states over one orbital period around Earth.
Figure 17: Local error in eccentric position states over one orbital period around Earth.
The Clohessy-Wiltshire (CW) equations, which are a popular form of linearized orbital mechanics deal with the relative motion between a chief and deputy spacecraft. As such, we compare this framework, modified to suit the relative orbital dynamics problem, in our work [68]. We invite the reader to read [68] for a more comprehensive analysis on how the Koopman operator can be used to relax some of the assumptions needed in the development of the CW equations. We present however, Figure 18 which highlights the ability of our Koopman framework to be quickly applied with minor modifications to this differing scenario. Presented is a comparison between our Koopman model, the CW equation solution and the nonlinear solution for a 200 second Monte Carlo simulation in which the chief spacecraft is in an elliptical orbit and the relative distance between both spacecraft is large, hence violating both assumptions of the CW equations. The true anomaly used in the CW equations, although incorrect, is an average true anomaly for the orbit and allows for a estimated comparison.
Figure 18: Comparison of the relative states between our Koopman framework, the CW equations and the nonlinear solution.
Plots of the error in the position state for both the Moon- and Jupiter-centered orbits allow us to analyze the effect of the size of the central body on the models ability to evolve the dynamics. A larger central body appears to have a greater periodic prediction error as compared to a central body that is smaller than that of the body the training data was derived from. As such, the possibility for the model to be used as a baseline, with additional data from the applied system added online, could enhance the performance of the model. With a short training time of around 120 min, it is not unreasonable to engage in the possibility of online data collection and learning to further enhance the accuracy of the predicted orbit.
@@ 5.2.2 Accuracy Metrics
We now look at the metrics provided in Section 4.2, and how the proposed Koopman operator is able to adhere to the conservative properties aforementioned. We see in Figure 19 that the values of all four metrics for the model around Earth are substantially low and very close to zero, hence illustrating the absolute error in each metric along the orbit is near zero. Furthermore, when these metrics are checked on orbits around both the Moon and Jupiter, over 10 orbital periods and with the J2 and SRP perturbations, Figures [20- 23], the conservation of the invariant properties is also preserved, as the values are also near zero.
The inclusion of the $r\cdot v$ loss function was specifically included because this metric has the highest degree of error among all other metrics. As seen with the 300 km Moon orbit, 5000 km orbit for 10 periods, it is possible that this metric be inconsistent with the nonlinear dynamics, however it was observed that after the inclusion of the specific loss function, the error in this metric was significantly reduced. As mentioned in Section 4.2, it is important to consider the magnitude of these values and whether averaging implies higher errors. These metrics prove the accuracy and reliability of our model in representing the 2BP not only insofar as its specific trained dataset is concerned, but also in its generalized applications to other systems.
@@
Figure 22: Invariant property metrics for the 2BP linearization over ten orbits.
Figure 23: Invariant property metrics for the J2 and SRP perturbed 2BP linearization.
@@ 5.2.3 Circular Restricted Three-Body Problem Linearization
Figure 25 shows the linearized evolution of the dynamics that is generated by our Koopman operator. It shows that the DNN is able to learn appropriate observables and lift the dimension of the states such that the K matrix was able to accurately predict the orbit of the CR3BP by rendering it as a linear system. Because the orbit, like the 2BP, is in-plane, the $z$-axis position and velocity are zero, therefore we are able to plot the propagation of the dynamics on a 2D plane to further inspect the accuracy of the propagation, as done in Figure 26. Note that in both Figures 25 and 26 the scale of the Earth and Moon representations are chosen for ease of viewing, and are not to scale with the propagated orbit. We also show the evolution of each of the 4 states independently in Figure 24, comparing them to the nonlinear dynamics, to emphasize the ability to closely follow the intricate dynamics of the CR3BP. It shall be noted that although the CR3BP is propagated in the non-dimensionalized units, in these figures the dimensionalized units are restored, and are displayed in the rotating frame.